# P. Guiraud, M. Hernández, A. Meyroneinc and A. Nogueira

### v1

@@ Abstract
We consider a parametric family of orientation-preserving piecewise contractions defined for every $x\in[0,1)$ by $f(x)=\lambda x+\delta+d\theta_{a}(x)\pmod{1}$, where $\lambda\in(0,1)$, $d\in(0,1-\lambda)$, $\delta\in[0,1]$, $a\in[0,1]$ and $\theta_{a}(x)=1$ if $x\geq a$ and $\theta_{a}(x)=0$ otherwise. In the special case where $a=1$, the family reduces to the well studied “contracted rotations” $x\mapsto\lambda x+\delta\pmod{1}$, which are 2-interval piecewise contractions when $\delta\in(1-\lambda,1)$. In the present paper, we allow any $a\in[0,1]$ in order to also consider maps with an additional discontinuity, that is, $3$-interval piecewise contractions. Supposing $\lambda$ and $d$ fixed, for any $\rho\in(0,1)$, we provide necessary and sufficient conditions on the values of the parameters $\delta$ and $a$ for $\rho$ to be the rotation number of the corresponding map. Under these conditions, we relate the map $f$ to a rotation $R_{\rho}$ of angle $\rho$, by giving the expression of a function $\phi:[0,1)\to\Lambda$ such that $f\circ\phi=\phi\circ R_{\rho}$, where $\Lambda$ is the support of the asymptotic dynamics of $f$. This enables us to study the dynamics as a function of $\delta$ and $a$ and to construct concrete examples of 3-interval piecewise contractions with some prescribed rotation number, attractor and symbolic dynamics.
@@ 1 Introduction
Let $\lambda\in(0,1)$, $d\in(0,1-\lambda)$ and $\delta,a\in[0,1]$. Consider the map $F:\mathbb{R}\to\mathbb{R}$ defined by
@@
where $\left\lfloor x\right\rfloor$ and $\{x\}$ are the integer part and the fractional part of $x$, respectively, and $\theta_{a}$ is the step function defined by
@@
We can verify that $F(x+1)=F(x)+1$ for all $x\in\mathbb{R}$ and that $F$ is monotonically increasing, since $d\in(0,1-\lambda)$. In other words, $F$ is a lift for the map $f:[0,1)\to[0,1)$ defined by
@@
As $F$ is non-decreasing, the quantity
$\rho_{f}:=\lim_{n\to\infty}\frac{F^{n}(x)}{n}\pmod{1},$
exists, does not depend on $x$ and is called the rotation number of $f$, see [3, 24].
In this paper, we are interested in the rotation number and, more generally, in the dynamical properties of $f$ as a function of the parameters $\delta$ and $a$. This kind of study has been done before for the commonly called contracted rotation, which is obtained by setting $d=0$ or $a=1$. In this case, and if $\delta\in(1-\lambda,1)$, the map $f$ is given for all $x\in[0,1)$ by
where $\eta=(1-\delta)/\lambda$. In particular, it has been shown [4, 5, 10, 16, 17] that the rotation number of this map is $\rho\in(0,1)$ if and only if
and $\delta(\rho^{-})$ is the left-sided limit of $\delta(\cdot)$ at $\rho$. Also, if (5) holds for some $\rho\in(0,1)$ the asymptotic dynamics of $f$ has the following properties. If $\rho\notin\mathbb{Q}$, then the closure of the attractor $\Lambda:=\bigcap_{n\in\mathbb{N}}f^{n}([0,1))$ of $f$ is a Cantor set. If $\rho\in\mathbb{Q}$ and $\delta\neq\delta(\rho)$, then $\Lambda$ is composed of a unique periodic orbit. Finally, the symbolic dynamics of $f$ on the set $\Lambda$, with respect to the natural partition $\{[0,\eta),[\eta,1)\}$, is the same as that of the rotation with respect to the partition $\{[0,1-\rho),[1-\rho,1)\}$.
As the map (4) is $\lambda$-Lipschitz with $\lambda\in(0,1)$ in each interval $[0,\eta)$ and $[\eta,1)$, it is a piecewise contraction (see [8], for instance). As it is also piecewise affine with the same slope $\lambda$ in each of the two intervals, it is called a 2-interval piecewise $\lambda$-affine contraction. The results on the family (4) have motivated recent studies on the qualitative dynamics of general interval piecewise contractions, with an arbitrary number of discontinuities and not necessarily piecewise affine. Among others, sharp bounds on the number of their periodic orbits [2, 7, 14, 20, 21], a decomposition of their attractor into periodic and Cantor sets components [7], and a characterization of the complexity of their symbolic dynamics [9, 23] have been provided during the last decade. However, the family (4) remains the only one for which there is some practical criterion, namely (5), to know the asymptotic and symbolic dynamics of its elements. On the other hand, it is a source of concrete examples, since (5) also allows constructing piecewise contractions with a specific asymptotic dynamics, e.g. with a Cantor set attractor, or a periodic orbit of a given period.
In this paper, we show that it is possible to extend the results and the method developed for (4), to maps with a larger number of discontinuities. As noted above, the lift of the map (4) is obtained by letting $a=1$ in (1), this way removing the discontinuities of $F$ that are not in $\mathbb{Z}$. The novelty of the present paper is to allow any $a\in[0,1]$ in order to give rise to projections (3) that are 3-interval generalizations of (4). For these maps, we give a necessary and sufficient condition on $\delta$ and $a$ that ensures that $f=f_{\delta,a}$ has a given rotation number $\rho$. It involves $\rho$, but also another number $\alpha\in[0,1]$ related to additional properties of $f_{\delta,a}$, concerning the relative position of its discontinuities, its number of periodic orbits and its symbolic dynamics. Then, we establish relations between the rigid rotations and the maps $f_{\delta,a}$ of the family (3). We give the expression of a function $\phi_{\delta,\rho,\alpha}:[0,1)\to\Lambda$ which maps the orbits of a rotation $R_{\rho}$ of angle $\rho$ to that of $f_{\delta,a}$ in its attractor, provided $\delta$ and $a$ are chosen such that $f_{\delta,a}$ has rotation number $\rho$. Under these conditions, the function $\phi_{\delta,\rho,\alpha}$ allows computing the points of $\Lambda$, thereby showing that its closure is a Cantor set for irrational $\rho$, and depending on $\alpha$ contains up to 2 periodic orbits for rational $\rho$. Also, the value of $\alpha$ selects, among the maps with rotation number $\rho$, those having in the set $\phi_{\delta,\rho,\alpha}([0,1))$ the same symbolic dynamics as $R_{\rho}$ with respect to a partition in three (or two) intervals given by $1-\rho$ and $\alpha$. Overall, we generalize the results obtained for (4), that allow constructing piecewise contractions with some prescribed rotation number, attractor and symbolic dynamics, to maps with an additional discontinuity.
@@
The results on the family (4) have been a starting point for further analyses, concerning the relations between the irrational nature of the rotation number of the map and the transcendental properties of the parameters $\lambda$ and $\delta$, or of the elements of its Cantor attractor [6, 17]. Additionally, the method developed for the map (4) has recently been extended to study 2-interval piecewise affine maps having a different slope in each interval [13, 18]. The map (4) has also found applications in the study of geodesics of homothety surfaces [1], voting methods [16], biological systems [11, 22], electronics [12] and finance [19]. This is why we hope that the present work may serve as a basis for further theoretical studies and will also be of interest for applied fields.
@@
The paper is organized as follows. In Section 2, we give the possible forms of the map (3) and present the main results of the paper. In Section 3, we show how the function $\phi_{\delta,\rho,\alpha}$, which relates (3) to a rotation, is constructed and we give its principal properties. In Sections 4, 5, 6 and 7 we provide the proofs of the main results. Finally, in Appendix A we gather some auxiliary and complementary results.
@@ 2.1 The different projections of $F$
which are the only two possible solutions of $F(x)=1$. Then, we obtain three different types for $f$ by letting $a\in[0,\eta_{1})$, $a\in[\eta_{1},\eta_{2}]$ and $a\in(\eta_{2},1]$. The first type exists if $\eta_{1}>0$ ($\delta<1-d$) and the third one if $\eta_{2}<1$ ($\delta>1-\lambda$). As $\eta_{1}<1$ and $0<\eta_{2}$, there is always some $a\in[0,1]\cap[\eta_{1},\eta_{2}]$ and the second type of map exists for any $\delta\in(1-\lambda-d,1)$. Now, we can verify that the three sets $M_{1},M_{2}$ and $M_{3}$ defined by
@@
As mentioned earlier, some maps given by the set $M$ can have a (ghost) fixed point. They are the maps such that $(\delta,a)\in\mathcal{F}_{1}\subset(M_{1}\cup M_{2})$, where
@@
$\mathcal{F}_{2}:=\left\{(\delta,a):\delta\in[1-d,1)\quad\text{and}\quad a\in\left[0,\frac{\delta+d-1}{1-\lambda}\right]\right\}\quad\text{with}\quad x_{*}=\frac{\delta+d-1}{1-\lambda},$ (11)
where $x_{*}$ denotes the corresponding fixed point. The sets $\mathcal{F}_{1}$ and $\mathcal{F}_{2}$ are shown in Figure 1. We say that $x_{*}$ is a ghost fixed point if $x_{*}=f(x^{-}_{*})\neq f(x_{*})$. Such a point exists only for the maps of $\mathcal{F}_{1}$ for which $x_{*}\in\{a,1\}$. A ghost fixed point is not in the image of the map, but the orbits accumulate on it and the map has zero rotation number. So, the sets $\mathcal{F}_{1}$ and $\mathcal{F}_{2}$ gather all the maps of $M$ with zero rotation number.
@@
Note that any orientation preserving 3-interval piecewise contraction defined on $[0,1)$ with constant positive slope and a gap of size 1 is of the form (7) or (8). Also, unless $a\in\{0,1\}$, when seen in the circle the maps defined by the set $M$ have two discontinuities, one at $x=0$ and the other one at $x=a$. However, as shown in Figure 2, the map (9) has an attracting invariant interval, namely $[f(a),f(a^{-}))$. Its restriction to this interval can be renormalized with an affine transformation to obtain the map (4), which has one discontinuity less. Also, the dynamical properties of this map can be deduced from those of (4). On the opposite, the maps (7) and (8) can be seen as renormalizations of some maps with gaps smaller than 1, and that are not necessarily orientation preserving, see Figure 3. So, (7) and (8) are also of interest to study some 3-interval piecewise contractions, with three discontinuities when seen in the circle.

### v3

@@ Abstract
We consider a family of piecewise contractions admitting a rotation number and defined for every $x\in[0,1)$ by $f(x)=\lambda x+\delta+d\theta_{a}(x)\pmod{1}$, where $\lambda\in(0,1)$, $d\in(0,1-\lambda)$, $\delta\in[0,1]$, $a\in[0,1]$ and $\theta_{a}(x)=1$ if $x\geq a$ and $\theta_{a}(x)=0$ otherwise. In the special case where $a=1$, the family reduces to the well studied “contracted rotations” $x\mapsto\lambda x+\delta\pmod{1}$, which are 2-interval piecewise $\lambda$-affine contractions when $\delta\in(1-\lambda,1)$. Considering $a\in(0,1)$ allows maps with an additional discontinuity, that is, $3$-interval piecewise $\lambda$-affine contractions. Supposing $\lambda$ and $d$ fixed, for any $\rho\in(0,1)$ and $\alpha\in[0,1]$, we provide the values of the parameters $\delta$ and $a$ for which the corresponding map has rotation number $\rho$, and a symbolic dynamics containing that of the rotation $R_{\rho}:[0,1)\to[0,1)$ of angle $\rho$ with respect to the partition given by the positions of $1-\rho$ and $\alpha$ in $[0,1)$. This enables in particular to determine the maps that have a given number of periodic orbits of an arbitrary period, or a Cantor set attractor supporting a dynamics of a given complexity.
@@ 1 Introduction
An interval map $f:I\to I$ is a piecewise contraction if there exist $\lambda\in(0,1)$ and a partition of $I$ in intervals $I_{1},I_{2},\dots,I_{N}$, such that $|f(x)-f(y)|\leq\lambda|x-y|$ for every $x,y\in I_{i}$ and $i\in\{1,\dots,N\}$. During the last decade, several results on the asymptotic dynamics of these maps have been obtained: we can mention the existence of a decomposition of their attractor into periodic and Cantor sets components [8], sharp bounds on the number of their periodic orbits [3, 8, 16, 23, 24], and the possible complexity functions of their symbolic dynamics [9, 26]. These results give the general characteristics of the asymptotic dynamics of interval piecewise contractions.
In theory, given a piecewise contraction, the number of periodic and Cantor set components of its attractor, as well as the complexity of its symbolic dynamics, can be estimated by analyzing the recurrence properties of the set of its discontinuities [8, 9]. In general, there is no explicit criterion relating directly the parameters of the map (such as the positions of the discontinuities or the contraction rate on each subinterval) to the exact structure and symbolic dynamics of its attractor. But when such a criterion can be established, it facilitates the detailed description of the dynamics as a function of the parameters, as well as the construction of examples whose attractor has some desired characteristics. It is the case for the 2-interval piecewise $\lambda$-affine contraction, often called contracted rotation, defined for any $x\in[0,1)$ by
where $\lambda\in(0,1)$, $\delta\in(1-\lambda,1)$ and $\eta=(1-\delta)/\lambda$. The following results, relating the parameters values of (1) to the properties of its attractor, have been proved in [5, 6, 10, 19, 20] by establishing explicit links between the rotations and (1). The map $f$ admits a rotation number (see (7)), that is equal to $\rho\in(0,1)$ if and only if
and $\delta(\rho^{-})$ is the left-sided limit of $\delta(\cdot)$ at $\rho$. When $\delta$ satisfies (2) for some $\rho\in(0,1)$ the asymptotic dynamics of $f$ has the following properties. If $\rho\notin\mathbb{Q}$, then the closure of the attractor $\Lambda:=\bigcap_{n\in\mathbb{N}}f^{n}([0,1))$ of $f$ is a Cantor set. If $\rho=p/q\in\mathbb{Q}$ and $\delta\neq\delta(\rho)$, then $\Lambda$ is a $q$-periodic orbit. In both cases, the symbolic dynamics of $f$ on the set $\Lambda$, with respect to the natural partition $\{[0,\eta),[\eta,1)\}$, is the same as that of the rotation $R_{\rho}$ of angle $\rho$ with respect to the partition $\{[0,1-\rho),[1-\rho,1)\}$. Besides, for irrational rotation number, the map restricted to the closure of $\Lambda$ is semi-conjugate to the rotation.
The map (1) has at most one periodic orbit and if its attractor is a Cantor set the code of any orbit on this set has the Sturmian complexity $p(n)=n+1$ of the irrational rotation. This illustrates the recent theoretical results mentioned earlier. On the one hand, injective piecewise contractions with N1N-1 discontinuities have at most $N$ periodic orbits [23]. If the map is right continuous and piecewise increasing, such as (1), its number of periodic orbits is bounded by NlN-l, where $l$ is the number of discontinuities whose image is equal to 0 (see [16]). On the other hand, for injective piecewise contractions, the complexity function of the code of any orbit (with respect to the natural partition) is an eventually affine function of slope at most N1N-1 [9, 26]. It follows that in order to observe a richer dynamics than that of (1) in the class of injective piecewise contractions, it is necessary to consider maps with more discontinuities.
It is known that the bounds on the number of periodic orbits and on the growth rate of the complexity functions are sharp. The sharpness of the first bound is proved by exhibiting a map with N1N-1 discontinuities and $N$ fixed points. The proof of the sharpness of the second one is harder, since maps with a non periodic attractor are exotic [14, 15, 16, 18, 19, 24, 27]. For instance, in [9] the map (1) is used as a base case to prove by induction the existence for any $N$ of maps with N1N-1 discontinuities and maximal complexity. Alternatively, in [26] the proof stands on showing that for any minimal interval exchange transformation there exists an injective piecewise contraction that is semi-conjugate to it. The family of the contracted rotations provides examples showing that the bounds are sharp for $N=2$. But for $N\geq 3$ arbitrary, examples with the maximal number of periodic orbits of any period or with a Cantor set attractor of maximal complexity remain difficult to construct. More generally, it is still missing a parametric family for which one can observe and control the features of the asymptotic dynamics arising when the number of discontinuities is greater than one.
The criterion (2) allows on the one hand a detailed description of the attractor of the map (1) as a function of the parameter $\delta$. On the other hand, by choosing $\rho$ adequately, (2) enables to construct examples having a stable periodic orbit of an arbitrary period, or a Cantor set attractor, in both cases with the symbolic dynamics of the corresponding rotation. In this paper, we show that it is possible to generalize methods developed for (1) in [10] to a larger family of maps having up to two discontinuities. For this family, we establish a criterion like (2), which relates directly the values of the parameters to some specific properties of the attractor. In particular, this enables to generate a wide set of examples with N1=2N-1=2 discontinuities for which the theoretical bounds on the number of periodic orbits and complexity are reached.
The family we consider is constructed as follows. We can verify that the map (1) admits a lift $F_{1}:\mathbb{R}\to\mathbb{R}$ defined by
$F_{1}(x)=\lambda x+\delta+(1-\lambda)\lfloor x\rfloor\quad\forall x\in\mathbb{R},$ (3)
where $\left\lfloor x\right\rfloor$ and $\{x\}$ are the integer part and the fractional part of $x$. In order to give rise to 3-interval generalizations of (1), we let $F:\mathbb{R}\to\mathbb{R}$ be defined by
@@
where $\lambda\in(0,1)$, $d\in(0,1-\lambda)$, $\delta,a\in[0,1]$ and $\theta_{a}$ is the step function defined by
@@
We are interested in the map defined by
@@
In the particular case where $a=1$ (or $d=0$), the map $f$ is given by (1). For $a\in(0,1)$ the map $F$ has 2 discontinuities in $[0,1)$. But depending on the value of $\delta\in(1-\lambda-d,1)$, as an interval map $f$ can have 1 or 2 discontinuities and therefore be a 2-interval or a 3-interval piecewise contraction (see Section 2 for an exhaustive study). We can verify that $F(x+1)=F(x)+1$ for all $x\in\mathbb{R}$ and that $F$ is a lift for $f$. Also, $f$ admits a rotation number
$\rho_{f}:=\lim_{n\to\infty}\frac{F^{n}(x)}{n}\pmod{1}$, (7)
since $F$ is monotonically increasing for $d\in(0,1-\lambda)$, which implies that (7) exists and does not depend on $x$, see [4, 28].
In this paper, for each $\lambda$ and $d$ fixed, we give the expression of subsets $\mathcal{P}_{\rho,\alpha}$ of $0,1]^{2}$, whose union over $\alpha\in[0,1]$ provides all the values of $(\delta,a)$ for which the rotation number of $f=f_{\delta,a}$ is $\rho\in(0,1)$. The number $\alpha$ allows to control additional features of $f_{\delta,a}$ such as its symbolic dynamics, its number of periodic orbits, and the relative position of $a$ with respect to the other discontinuity. Indeed, we show that when $(\delta,a)$ belongs to $\mathcal{P}_{\rho,\alpha}$, the relation $f_{\delta,a}\circ\phi_{\delta,\rho,\alpha}=\phi_{\delta,\rho,\alpha}\circ R_{\rho}$ holds for an explicit function $\phi_{\delta,\rho,\alpha}:[0,1)\to\Lambda=\bigcap_{n\in\mathbb{N}}f_{\delta,a}^{n}([0,1))$ which is right continuous and non-decreasing. It allows to prove that the symbolic dynamics of $f_{\delta,a}$ contains that of the rotation $R_{\rho}$, with respect to the partition in three (or two) intervals given by $1-\rho$ and $\alpha$. If $\{\alpha\}$ does not belong to the orbit (backward and forward) of $1-\rho$ by $R_{\rho}$, then $\Lambda$ is the image of $\phi_{\delta,\rho,\alpha}$ and either it is the union of 2 periodic orbits ($\rho$ rational) or its closure is a Cantor set ($\rho$ irrational) supporting a dynamics of complexity $p(n)=2n+1$. Otherwise, $\Lambda$ supports a dynamics of any eventually affine complexity (of slope 0 or 1) determined by $\rho$ and the position of $\{\alpha\}$ in the orbit of $1-\rho$.
As a consequence, we provide the values of the parameters $\delta$ and $a$ for which the maps (6) have a given number of periodic orbits of an arbitrary period or a Cantor set attractor with a given affine complexity: to obtain a map $f_{\delta,a}$ whose attractor has some desired properties, it is enough to choose $\rho$ and $\alpha$ accordingly and to pick $(\delta,a)$ in the set $\mathcal{P}_{\rho,\alpha}$. The rotation number $\rho$ alone determines only the Cantor set or periodic nature of the attractor. Together, $\rho$ and $\alpha$ set the symbolic dynamics, that in turn determines the number of periodic orbits and the complexity.
We trust that our results can be generalized to $N$-interval piecewise $\lambda$-affine contractions, provided they admit an increasing lift. It would give the parameters of the maps having any prescribed rotation number, number of periodic orbits and complexity, within the admissible values of these quantities in such a family. Also, every injective piecewise contraction of $N$ intervals is topologically conjugate to a piecewise affine contraction of $N$ intervals whose slopes are 1/2 in absolute value [23]. Therefore, an extension of our work to $N$-interval piecewise $1/2$-affine contractions would also be of interest for the study of general order-preserving injective piecewise contractions as initiated in [3]. The approach used in the present paper consists in establishing explicit relations between some interval exchange transformations, namely the rotations, and the piecewise contractions (6). The study of injective maps with some decreasing branches, may require to consider more general interval exchange transformations, such as those with flips, as it is the case in [13].
@@
The results on (1) have been a starting point for further analysis concerning the relation between the algebraic properties of the parameters $\lambda$ and $\delta$ and the rational nature of the rotation number of the map [20]. Also, the transcendental nature of the elements of the Cantor attractors of the maps (1) having as slope the reciprocal of an integer has been established [7]. On the other hand, the map (1) appears in the study of geodesics of homothety surfaces [2], period adding bifurcations [17], voting methods [19], biological systems [11, 25], electronics [12] and finance [22]. This is why we think that the present work, that extends the framework and results on (1), can serve as a basis for further theoretical studies and will also be of interest for applied fields.
@@
The paper is organized as follows. In Section 2, we give the possible forms of the map (6) and present the main results of the paper. In Section 3, we show how to construct the function $\phi_{\delta,\rho,\alpha}$ which relates a rotation to maps of the family (6) and we give its principal properties. In Sections 4, 5, 6 and 7 we provide the proofs of the main results. Finally, in Appendix A we gather some auxiliary and complementary materials.
@@ 2.1 The different projections of $F$
which are the only two possible solutions of $F(x)=1$. Then, we obtain three different types for $f$ by letting $a\in[0,\eta_{1})$, $a\in[\eta_{1},\eta_{2}]$ or $a\in(\eta_{2},1]$. The first type exists if $\eta_{1}>0$ ($\delta<1-d$) and the third one if $\eta_{2}<1$ ($\delta>1-\lambda$). As $\eta_{1}<1$ and $0<\eta_{2}$, there is always some $a\in[0,1]\cap[\eta_{1},\eta_{2}]$ and the second type of map exists for any $\delta\in(1-\lambda-d,1)$. Now, we can verify that the three sets $M_{1},M_{2}$ and $M_{3}$ defined by
@@
As mentioned earlier, some maps given by the set $M$ can have a (ghost) fixed point $x_{*}$. They are the maps such that $(\delta,a)\in\mathcal{F}_{1}\subset(M_{1}\cup M_{2})$, where
@@
$\mathcal{F}_{2}:=\left\{(\delta,a):\delta\in[1-d,1)\quad\text{and}\quad a\in\left[0,\frac{\delta+d-1}{1-\lambda}\right]\right\}\quad\text{with}\quad x_{*}=\frac{\delta+d-1}{1-\lambda}.$ (13)
The sets $\mathcal{F}_{1}$ and $\mathcal{F}_{2}$ are shown in Figure 1. We say that $x_{*}$ is a ghost fixed point if $x_{*}=f(x^{-}_{*})$ but $x_{*}\neq f(x_{*})$ or $x_{*}\notin f([0,1))$. Such a point exists only for the maps of $\mathcal{F}_{1}$ for which $x_{*}\in\{a,1\}$. In our case, a ghost fixed point is not in the image of the map, but the orbits accumulate on it and the map has zero rotation number. So, the sets $\mathcal{F}_{1}$ and $\mathcal{F}_{2}$ gather all the maps of $M$ with zero rotation number.
@@
Note that, unless $a\in\{0,1\}$, when seen in the circle the maps defined by the set $M$ have two discontinuities, one at $x=0$ and the other one at $x=a$. However, as shown in Figure 2, the map (11) has an attracting invariant interval, namely $f(a),f(a^{-}))$. Its restriction to this interval can be renormalized with an affine transformation to obtain the map (1), which has one discontinuity less. The dynamical properties of this map can be deduced from those of (1). On the opposite, the maps (9) and (10) can be seen as renormalizations of some maps with a discontinuity jump smaller than 1, and that are not necessarily orientation preserving, see Figure 3. So, (9) and (10) are also of interest to study some 3-interval piecewise contractions, with a discontinuity jump smaller than 1 and three discontinuities when seen in the circle.